# Counting and Entropy Bounds for Structure-Avoiding Spatially-Coupled LDPC Constructions

### v1

@@ IV Conclusion
By casting structure-avoiding QC-SC-LDPC design as a CSP, we derived explicit lower bounds on the number of feasible edge-spreading and lifting assignments via a quantitative CLLL, allowing general non-uniform spreading and lifting distributions. We further obtained a computable diversity lower bound on the number of distinct solutions producible by the MT resampling algorithm using Rényi-entropy bounds, with closed-form specializations for eliminating $4$-cycle candidates. These bounds quantify the feasible design space and provide practical guidance for choosing memory/lifting parameters and sizing the remaining code-search space.
@@ Appendix A Proof of the Theorem 3
$\Big(\sum_{s\in S}D(s)\Big)^{\alpha}\leq|S|^{\alpha-1}\sum_{s\in S}D(s)^{\alpha}.$
@@
$\sum_{s\in S}D(s)^{\alpha}\geq|S|^{1-\alpha}.$
Taking logarithms and multiplying by $\frac{1}{1-\alpha}$ yields (A.1).
Applying (A.1) to $D=D_{MT}$ gives
$H_{\alpha}(D_{MT})\leq\ln|S|\quad\Longrightarrow\quad|S|\geq\exp\big(H_{\alpha}(D_{MT})\big).$
@@
$_{MT}\Big(\bigcap_{i\in[k]}\overline{H_{i}}\Big)=|S|\geq\exp\big(H_{\alpha}(D_{MT})\big),$
@@
Appendix B Proof of the Corollary 3
@@
$=\frac{\gamma\kappa}{1-\alpha}\ln\Bigg(\sum_{\ell=0}^{m_{t}}p_{\ell}^{\alpha}\Bigg)+\frac{\gamma\kappa}{1-\alpha}\ln\Bigg(\sum_{t\in\mathbb{Z}_{Z}}q_{t}^{\alpha}\Bigg).$ (B.2)
$H_{\alpha}(D_{MT})\geq H_{\alpha}(\Omega)-\frac{\alpha}{\alpha-1}ln\sum_{I\in Ind(\mathcal{A})}\prod_{A\in I}\mu(A)$ (B.3)
$\geq\frac{1}{1-\alpha}ln\sum_{t\in\mathcal{S}}P_{\mathcal{T}}(t)^{\alpha}-\frac{\alpha}{\alpha-1}\sum_{A\in\mathcal{A}}\mu(A)$ (B.4)
$=\frac{1}{1-\alpha}ln\sum_{t\in\mathcal{S}}[Z(m_{t}+1)]^{-\gamma\kappa\alpha}-\frac{\alpha}{\alpha-1}\sum_{A\in\mathcal{A}}\frac{1}{\Delta-1}$ (B.5)
$=\frac{1}{1-\alpha}ln[Z(m_{t}+1)]^{(1-\alpha)\gamma\kappa}-\frac{\alpha}{\alpha-1}\frac{\binom{\gamma}{2}\binom{\kappa}{2}}{\Delta-1}$ (B.6)
$\overset{\alpha\to\infty}{=}ln[Z(m_{t}+1)]^{\gamma\kappa}-\frac{\binom{\gamma}{2}\binom{\kappa}{2}}{\Delta-1}$ (B.7)
$\#_{MT}(\bigcap_{i\in[n]}\bar{A_{i}})$ $\geq exp(H_{\alpha}(D_{MT}))$ (B.8)
$\geq\frac{[Z(m_{t}+1)]^{\gamma\kappa}}{exp(\frac{\binom{\gamma}{2}\binom{\kappa}{2}}{4\gamma\kappa-6\gamma-6\kappa+8})}$ (B.9)
@@
$H_{\alpha}(D_{MT})\geq H_{\alpha}(\Omega)-\frac{\alpha}{\alpha-1}ln\sum_{I\in Ind(\mathcal{A})}\prod_{A\in I}\mu(A)$ (B.10)
$\geq\frac{1}{1-\alpha}ln\sum_{t\in\mathcal{S}}P_{\mathcal{T}}(t)^{\alpha}-\frac{\alpha}{\alpha-1}ln\prod_{i\in[n]}(1+\sum_{\begin{subarray}{c}A\in\mathcal{A}\\ i\in var(A)\end{subarray}}l(A))$ (B.11)
For $\mathcal{H}=\mathcal{C}_{4}$, $|\mathrm{var}(A)|=4$
$l(A)$ $=(1+\mu(A))^{\frac{1}{|var(A)|}}-1=\left(\frac{\Delta}{\Delta-1}\right)^{\frac{1}{4}}-1$ (B.12)
$=\left(\frac{(2\gamma-3)(2\kappa-3)}{(2\gamma-3)(2\kappa-3)-1}\right)^{\frac{1}{4}}-1$ (B.13)
$ln\prod_{i\in[n]}(1+\sum_{\begin{subarray}{c}A\in\mathcal{A}\\ i\in var(A)\end{subarray}}l(A))=ln\prod_{i\in[n]}(1+Wl(A))$ (B.14)
$=(\gamma\kappa)ln\left(1+(\gamma-1)(\kappa-1)l(A)\right)$ (B.15)
@@
$H_{\alpha}(D_{MT})$ (B.16)
$\geq\frac{1}{1-\alpha}ln\sum_{t\in\mathcal{S}}P_{\mathcal{T}}(t)^{\alpha}-\frac{\alpha}{\alpha-1}ln\prod_{i\in[n]}(1+\sum_{\begin{subarray}{c}A\in\mathcal{A}\\ i\in var(A)\end{subarray}}l(A))$ (B.17)
$=\frac{1}{1-\alpha}ln\sum_{t\in\mathcal{S}}[Z(m_{t}+1)]^{-\gamma\kappa\alpha}$ (B.18)
$-\frac{\alpha}{\alpha-1}(\gamma\kappa)ln\left(1+(\gamma-1)(\kappa-1)l(A)\right)$ (B.19)
$\overset{\alpha\to\infty}{=}ln\left(\frac{Z(m_{t}+1)}{\left(1+(\gamma-1)(\kappa-1)l(A)\right)}\right)^{\gamma\kappa}$ (B.20)
$\#_{MT}(\bigcap_{i\in[n]}\bar{A_{i}})$ $\geq exp(H_{\alpha}(D_{MT}))$ (B.21)
$\geq\left(\frac{Z(m_{t}+1)}{\left(1+(\gamma-1)(\kappa-1)l(A)\right)}\right)^{\gamma\kappa}$ (B.22)
Appendix C Proof of the Comparison Inequality
@@
$X$ $:=Z(m_{t}+1),$ (C.1)
$D$ $:=4\gamma\kappa-6\gamma-6\kappa+8,$ (C.2)
$l(A)$ $:=\left(\frac{(2\gamma-3)(2\kappa-3)}{(2\gamma-3)(2\kappa-3)-1}\right)^{\frac{1}{8}}-1,$ (C.3)
$u$ $:=(\gamma-1)(\kappa-1)l(A),$ (C.4)
$E$ $:=\frac{\binom{\gamma}{2}\binom{\kappa}{2}}{D}.$ (C.5)
@@
$\left(\frac{X}{1+u}\right)^{\gamma\kappa}\;>\;\frac{X^{\gamma\kappa}}{\exp(E)}.$ (C.6)
@@
$(2\gamma-3)(2\kappa-3)=4\gamma\kappa-6\gamma-6\kappa+9=D+1.$ (C.7)
@@
$D=(2\gamma-3)(2\kappa-3)-1\geq 24>0.$ (C.8)
Combining (C.3) and (C.7) yields
$l(A)=\left(1+\frac{1}{D}\right)^{\frac{1}{8}}-1>0,$ (C.9)
@@
$f^{\prime\prime}(x)=\frac{1}{8}\Bigl(\frac{1}{8}-1\Bigr)(1+x)^{\frac{1}{8}-2}<0.$ (C.10)
@@
$(1+x)^{\frac{1}{8}}<1+\frac{x}{8},\qquad x>0.$ (C.11)
Applying (C.11) with $x=1/D$ gives
$l(A)=\left(1+\frac{1}{D}\right)^{\frac{1}{8}}-1<\frac{1}{8D},$ (C.12)
@@
$u=(\gamma-1)(\kappa-1)l(A)<\frac{(\gamma-1)(\kappa-1)}{8D}.$ (C.13)
@@
$\ln(1+y)<y,$ (C.14)
because $g(y)=y-\ln(1+y)$ satisfies $g(0)=0$ and $g^{\prime}(y)=y/(1+y)>0$ for $y>0$. Using (C.14) with $y=u>0$ and exponentiating yields
$(1+u)^{\gamma\kappa}=\exp\!\bigl(\gamma\kappa\ln(1+u)\bigr)<\exp(\gamma\kappa u).$ (C.15)
@@
$E=\frac{\binom{\gamma}{2}\binom{\kappa}{2}}{D}=\frac{\gamma\kappa(\gamma-1)(\kappa-1)}{4D}.$ (C.16)
Combining (C.13) and (C.16) we obtain
$\gamma\kappa u<\gamma\kappa\cdot\frac{(\gamma-1)(\kappa-1)}{8D}=\frac{E}{2}<E,$ (C.17)
@@
$(1+u)^{\gamma\kappa}<\exp(\gamma\kappa u)<\exp(E).$ (C.18)
Taking reciprocals in (C.18) (all terms are positive) and multiplying by $X^{\gamma\kappa}>0$ proves (C.6). ∎
Appendix D Numerical Verification for $(\gamma,\kappa)=(3,5)$
Although $\gamma=3$ does not satisfy $\gamma>3$, the comparison in (C.6) still holds whenever $D>0$. For $(\gamma,\kappa)=(3,5)$ we have
$D$ $=4\gamma\kappa-6\gamma-6\kappa+8=4\cdot 15-18-30+8=20,$ (D.1)
$l(A)$ $=\left(1+\frac{1}{D}\right)^{\frac{1}{8}}-1=\left(\frac{21}{20}\right)^{\frac{1}{8}}-1$ (D.2)
$\approx 0.00611740588712562,$ (D.3)
$u$ $=(\gamma-1)(\kappa-1)l(A)=8l(A)$ (D.4)
$\approx 0.0489392470970049,$ (D.5)
$E$ $=\frac{\binom{3}{2}\binom{5}{2}}{D}=\frac{3\cdot 10}{20}=1.5$ (D.6)
$\exp(E)$ $\approx 4.48168907033806.$ (D.7)
@@ Appendix D Numerical Verification for $(\gamma,\kappa)=(3,5)$
$(1+u)^{\gamma\kappa}$ $=(1+u)^{15}\approx(1.0489392470970049)^{15}$ (D.8)
$\approx 2.04764671546561$ (D.9)
$<\exp(E)\approx 4.48168907033806.$ (D.10)
@@
$\left(\frac{1}{1+u}\right)^{15}$ $\approx 0.488365494129007$ (D.11)
$>\exp(-E)=e^{-1.5}\approx 0.223130160148430,$ (D.12)

### v2

@@ IV Conclusion
By casting structure-avoiding QC-SC-LDPC design as a CSP, we derived explicit lower bounds on the number of feasible edge-spreading and lifting assignments via a quantitative CLLL, allowing general non-uniform spreading and lifting distributions. We further obtained a computable diversity lower bound on the number of distinct solutions producible by the MT resampling algorithm using Rényi entropy bounds, with closed-form specializations for eliminating $4$-cycles. These bounds quantify the feasible design space and provide practical guidance for choosing memory/lifting parameters and sizing the remaining code-search space.
@@ Appendix A Proof of the Theorem 1
Let $\Omega:=\big(\mathbf{a}\times\mathbb{Z}_{Z}\big)^{\gamma\kappa}=[(a_{0},a_{1},\ldots,a_{m_{t}})\times(0,1,\ldots,Z-1)]^{\gamma\kappa}$ be the assignment space of all edge-variables $X_{i,j}=(\mathbf{P}(i,j),\mathbf{L}(i,j))$, and let
$S:=\bigcap_{i\in[k]}\bar{H_{i}}.$ (A.1)
$\Pr_{\Omega}(\omega)$ $=\prod_{i\in[\gamma],\,j\in[\kappa]}\Pr[\mathbf{P}(i,j)=\omega^{P}_{i,j}]\Pr[\mathbf{L}(i,j)=\omega^{L}_{i,j}]$
$\leq(p_{\max}q_{\max})^{\gamma\kappa}.$
Hence,
$\Pr_{\Omega}(S)=\sum_{\omega\in S}\Pr_{\Omega}(\omega)\leq\#(S)\cdot(p_{\max}q_{\max})^{\gamma\kappa}.$ (A.2)
which implies
$(S)\geq\left(\frac{1}{p_{\max}q_{\max}}\right)^{\gamma\kappa}\Pr_{\Omega}(S).$ (A.3)
By the quantitative clique-LLL bound used in [9, Lemma 2] (under the stated condition on $\max_{i}P_{H_{i}}$),
$\Pr_{\Omega}(S)\geq\prod_{v\in[m]}\Big(1-\sum_{i\in K_{v}}x_{i,v}\Big).$ (A.4)
Combining with (A.3) yields
$(S)\geq\left(\frac{1}{p_{\max}q_{\max}}\right)^{\gamma\kappa}\prod_{v\in[m]}\Big(1-\sum_{i\in K_{v}}x_{i,v}\Big).$ (A.5)
Finally, substituting into [9, Theorem 1] the same choice of $\{x_{i,v}\}$ that yields the two explicit cases II and IIII gives the claimed bound.
Appendix B Proof of the Theorem 3
Since Lemmas 3, 4, and 5 are established in  [8] based on the Quantitative Cluster Expansion Lovász Local Lemma (Lemma 7), we adopt this specific formulation in the remainder of this paper for the sake of convenience. We note that these results can be readily generalized to more general forms of the Lovász Local Lemma; such extensions will be provided in an extended version of this work.In the context of the Quantitative Cluster Expansion Lovász Local Lemma, we can define the output distribution $D_{MT}$ of the MT algorithm analogously to the approach taken in [9].
Definition 5
The distribution of $\Omega$ conditioned on avoiding $\mathcal{A}$ is called the LLL-distribution. The distribution at the M-T algorithm termination is called the MT-distribution.
@@ Appendix B Proof of the Theorem 3
$\Big(\sum_{s\in S}D(s)\Big)^{\alpha}\leq|S|^{\alpha-1}\sum_{s\in S}D(s)^{\alpha}.$ (B.2)
@@
$\sum_{s\in S}D(s)^{\alpha}\geq|S|^{1-\alpha}.$ (B.3)
Taking logarithms and multiplying by $\frac{1}{1-\alpha}$ yields (B.1).
Applying (B.1) to $D=D_{MT}$ gives
$H_{\alpha}(D_{MT})\leq\ln|S|\quad\Longrightarrow\quad|S|\geq\exp\big(H_{\alpha}(D_{MT})\big).$ (B.4)
@@
$_{MT}\Big(\bigcap_{i\in[k]}\overline{H_{i}}\Big)=|S|\geq\exp\big(H_{\alpha}(D_{MT})\big),$ (B.5)
@@
Appendix C Proof of the Corollary 3
Motivated by the reasons outlined in the proof of Theorem 3, and to facilitate the construction of a simple and concrete example, we focus on the Quantitative Lovász Local Lemma 6. Note that for a given dependency graph $G_{D}$, the inclusion $\mathcal{I}_{a}(G_{D})\subset\mathcal{I}_{CE}(G_{D})$ holds ( Let $\mathcal{I}_{a}(G_{D})$ and $\mathcal{I}_{CE}(G_{D})$ be the interiors corresponding to Lemma 6 and Lemma 7, respectively; for details, refer to Appendix F). Consequently, whenever the conditions of the Quantitative Lovász Local Lemma in Corollary 3 are satisfied, the conditions of Theorem 3 are automatically met. Following a line of reasoning similar to that in [21], it follows that when (18) holds:
@@ Appendix C Proof of the Corollary 3
$=\frac{\gamma\kappa}{1-\alpha}\ln\Bigg(\sum_{\ell=0}^{m_{t}}p_{\ell}^{\alpha}\Bigg)+\frac{\gamma\kappa}{1-\alpha}\ln\Bigg(\sum_{t\in\mathbb{Z}_{Z}}q_{t}^{\alpha}\Bigg).$ (C.1)
$H_{\alpha}(D_{MT})\geq H_{\alpha}(\Omega)-\frac{\alpha}{\alpha-1}ln\sum_{I\in Ind(\mathcal{A})}\prod_{A\in I}\mu(A)$ (C.2)
$\geq\frac{1}{1-\alpha}ln\sum_{t\in\mathcal{S}}P_{\mathcal{T}}(t)^{\alpha}-\frac{\alpha}{\alpha-1}\sum_{A\in\mathcal{A}}\mu(A)$ (C.3)
$=\frac{1}{1-\alpha}ln\sum_{t\in\mathcal{S}}[Z(m_{t}+1)]^{-\gamma\kappa\alpha}-\frac{\alpha}{\alpha-1}\sum_{A\in\mathcal{A}}\frac{1}{\Delta-1}$ (C.4)
$=\frac{1}{1-\alpha}ln[Z(m_{t}+1)]^{(1-\alpha)\gamma\kappa}-\frac{\alpha}{\alpha-1}\frac{\binom{\gamma}{2}\binom{\kappa}{2}}{\Delta-1}$ (C.5)
$\overset{\alpha\to\infty}{=}ln[Z(m_{t}+1)]^{\gamma\kappa}-\frac{\binom{\gamma}{2}\binom{\kappa}{2}}{\Delta-1}.$ (C.6)
$\#_{MT}(\bigcap_{i\in[n]}\bar{A_{i}})$ $\geq exp(H_{\alpha}(D_{MT}))$ (C.7)
$\geq\frac{[Z(m_{t}+1)]^{\gamma\kappa}}{exp(\frac{\binom{\gamma}{2}\binom{\kappa}{2}}{4\gamma\kappa-6\gamma-6\kappa+8})}.$ (C.8)
@@
$H_{\alpha}(D_{MT})\geq H_{\alpha}(\Omega)-\frac{\alpha}{\alpha-1}ln\sum_{I\in Ind(\mathcal{A})}\prod_{A\in I}\mu(A)$ (C.9)
$\geq\frac{1}{1-\alpha}ln\sum_{t\in\mathcal{S}}P_{\mathcal{T}}(t)^{\alpha}-\frac{\alpha}{\alpha-1}ln\prod_{i\in[n]}(1+\sum_{\begin{subarray}{c}A\in\mathcal{A}\\ i\in var(A)\end{subarray}}l(A)).$ (C.10)
For $\mathcal{H}=\mathcal{C}_{4}$, $|\mathrm{var}(A)|=8$
$l(A)$ $=(1+\mu(A))^{\frac{1}{|var(A)|}}-1=\left(\frac{\Delta}{\Delta-1}\right)^{\frac{1}{8}}-1$ (C.11)
$=\left(\frac{(2\gamma-3)(2\kappa-3)}{(2\gamma-3)(2\kappa-3)-1}\right)^{\frac{1}{8}}-1.$ (C.12)
$ln\prod_{i\in[n]}(1+\sum_{\begin{subarray}{c}A\in\mathcal{A}\\ i\in var(A)\end{subarray}}l(A))=ln\prod_{i\in[n]}(1+Wl(A))$ (C.13)
$=(\gamma\kappa)ln\left(1+(\gamma-1)(\kappa-1)l(A)\right).$ (C.14)
@@
$H_{\alpha}(D_{MT})$ (C.15)
$\geq\frac{1}{1-\alpha}ln\sum_{t\in\mathcal{S}}P_{\mathcal{T}}(t)^{\alpha}-\frac{\alpha}{\alpha-1}ln\prod_{i\in[n]}(1+\sum_{\begin{subarray}{c}A\in\mathcal{A}\\ i\in var(A)\end{subarray}}l(A))$ (C.16)
$=\frac{1}{1-\alpha}ln\sum_{t\in\mathcal{S}}[Z(m_{t}+1)]^{-\gamma\kappa\alpha}$ (C.17)
$-\frac{\alpha}{\alpha-1}(\gamma\kappa)ln\left(1+(\gamma-1)(\kappa-1)l(A)\right)$ (C.18)
$\overset{\alpha\to\infty}{=}ln\left(\frac{Z(m_{t}+1)}{\left(1+(\gamma-1)(\kappa-1)l(A)\right)}\right)^{\gamma\kappa}.$ (C.19)
$\#_{MT}(\bigcap_{i\in[n]}\bar{A_{i}})$ $\geq exp(H_{\alpha}(D_{MT}))$ (C.20)
$\geq\left(\frac{Z(m_{t}+1)}{\left(1+(\gamma-1)(\kappa-1)l(A)\right)}\right)^{\gamma\kappa}.$ (C.21)
Appendix D Proof of the Comparison Inequality
@@
$X$ $:=Z(m_{t}+1),$ (D.1)
$D$ $:=4\gamma\kappa-6\gamma-6\kappa+8,$ (D.2)
$l(A)$ $:=\left(\frac{(2\gamma-3)(2\kappa-3)}{(2\gamma-3)(2\kappa-3)-1}\right)^{\frac{1}{8}}-1,$ (D.3)
$u$ $:=(\gamma-1)(\kappa-1)l(A),$ (D.4)
$E$ $:=\frac{\binom{\gamma}{2}\binom{\kappa}{2}}{D}.$ (D.5)
@@
$\left(\frac{X}{1+u}\right)^{\gamma\kappa}\;>\;\frac{X^{\gamma\kappa}}{\exp(E)}.$ (D.6)
@@
$(2\gamma-3)(2\kappa-3)=4\gamma\kappa-6\gamma-6\kappa+9=D+1.$ (D.7)
@@
$D=(2\gamma-3)(2\kappa-3)-1\geq 24>0.$ (D.8)
Combining (D.3) and (D.7) yields
$l(A)=\left(1+\frac{1}{D}\right)^{\frac{1}{8}}-1>0,$ (D.9)
@@
$f^{\prime\prime}(x)=\frac{1}{8}\Bigl(\frac{1}{8}-1\Bigr)(1+x)^{\frac{1}{8}-2}<0.$ (D.10)
@@
$(1+x)^{\frac{1}{8}}<1+\frac{x}{8},\qquad x>0.$ (D.11)
Applying (D.11) with $x=1/D$ gives
$l(A)=\left(1+\frac{1}{D}\right)^{\frac{1}{8}}-1<\frac{1}{8D},$ (D.12)
@@
$u=(\gamma-1)(\kappa-1)l(A)<\frac{(\gamma-1)(\kappa-1)}{8D}.$ (D.13)
@@
$\ln(1+y)<y,$ (D.14)
because $g(y)=y-\ln(1+y)$ satisfies $g(0)=0$ and $g^{\prime}(y)=y/(1+y)>0$ for $y>0$. Using (D.14) with $y=u>0$ and exponentiating yields
$(1+u)^{\gamma\kappa}=\exp\!\bigl(\gamma\kappa\ln(1+u)\bigr)<\exp(\gamma\kappa u).$ (D.15)
@@
$E=\frac{\binom{\gamma}{2}\binom{\kappa}{2}}{D}=\frac{\gamma\kappa(\gamma-1)(\kappa-1)}{4D}.$ (D.16)
Combining (D.13) and (D.16) we obtain
$\gamma\kappa u<\gamma\kappa\cdot\frac{(\gamma-1)(\kappa-1)}{8D}=\frac{E}{2}<E,$ (D.17)
@@
$(1+u)^{\gamma\kappa}<\exp(\gamma\kappa u)<\exp(E).$ (D.18)
Taking reciprocals in (D.18) (all terms are positive) and multiplying by $X^{\gamma\kappa}>0$ proves (D.6). ∎
Appendix E Numerical Verification for $(\gamma,\kappa)=(3,5)$
For $(\gamma,\kappa)=(3,5)$ we have
$D$ $=4\gamma\kappa-6\gamma-6\kappa+8=4\cdot 15-18-30+8=20,$ (E.1)
$l(A)$ $=\left(1+\frac{1}{D}\right)^{\frac{1}{8}}-1=\left(\frac{21}{20}\right)^{\frac{1}{8}}-1$ (E.2)
$\approx 0.00611740588712562,$ (E.3)
$u$ $=(\gamma-1)(\kappa-1)l(A)=8l(A)$ (E.4)
$\approx 0.0489392470970049,$ (E.5)
$E$ $=\frac{\binom{3}{2}\binom{5}{2}}{D}=\frac{3\cdot 10}{20}=1.5$ (E.6)
$\exp(E)$ $\approx 4.48168907033806.$ (E.7)
@@ Appendix E Numerical Verification for $(\gamma,\kappa)=(3,5)$
$(1+u)^{\gamma\kappa}$ $=(1+u)^{15}\approx(1.0489392470970049)^{15}$ (E.8)
$\approx 2.04764671546561$ (E.9)
$<\exp(E)\approx 4.48168907033806.$ (E.10)
@@
$\left(\frac{1}{1+u}\right)^{15}$ $\approx 0.488365494129007$ (E.11)
$>\exp(-E)=e^{-1.5}\approx 0.223130160148430,$ (E.12)
@@
Appendix F Auxiliary Lemmas
The Lovász local lemma (LLL), introduced by Erdős and Lovász in 1975, has become one of the most important probabilistic methods. LLL provides sufficient conditions under which a set of undesirable events $\mathcal{A}$ in a probability space $\Omega$ can be simultaneously avoided, meaning the conditions for $\Pr(\cap_{A\in\mathcal{A}}\bar{A})>0$ to hold. The relationships between these bad events are depicted by an undirected graph called the dependency graph $G_{D}=([n],E)$, where each bad event corresponds to a vertex in the dependency graph. An event $A_{i}$ is independent of $\{A_{j}:j\neq i,j\notin\mathcal{N}(i)\}$, where $\mathcal{N}(i)$ stands for the neighborhood of $A_{i}$ in $G$.
Note that in the variable-generated system model, the event-variable graph $G_{B}=([n],[m^{\prime}],E)$ not only captures the dependency relationships between events but also the strength of these dependencies, which the dependency graph $G_{D}=([n],E)$ cannot capture. Therefore, the event-variable graph contains more information than the dependency graph. Given the bipartite graph $G_{B}=([n],[m^{\prime}],E)$, let $E^{\prime}=\{(v_{1},v_{2}):v_{1},v_{2}\in[n],\mathcal{N}(v_{1})\cap\mathcal{N}(v_{2})\neq\emptyset\}$. Define the dependency graph corresponding to the event-variable graph $G_{B}=([n],[m^{\prime}],E)$ as $G_{D}(G_{B})=([n],E^{\prime})$.
We use $\mathcal{A}\sim(G_{D},\mathbf{p})$ to denote that (i) $G_{D}$ is a dependency graph of $\mathcal{A}$ and (ii) the probability vector of $\mathcal{A}$ is $\mathbf{p}$. Given a dependency graph $G_{D}$, define the abstract interior $\mathcal{I}_{a}(G_{D})$ to be the set consisting of all vectors $\mathbf{p}$ such that $\Pr\left(\bigcap_{A\in\mathcal{A}}A\right)>0$ for any $\mathcal{A}\sim(G_{D},\mathbf{p})$. We abbreviate $\mathcal{I}{a}(G_{D}(G_{B}))$ as $\mathcal{I}{a}(G_{B})$. Similarly, We use $\mathcal{A}\sim(G_{B},\mathbf{p})$ to denote that (i) $H_{E}$ is a event-variable graph of $\mathcal{A}$ and (ii) the probability vector of $\mathcal{A}$ is $\mathbf{p}$. In this appendix, we reuse the symbol $\mathbf{p}$ to denote a set of probability vectors, distinct from its usage in the main text.
LLL offers a powerful method to demonstrate the existence of complex combinatorial objects that satisfy a given set of requirements. The initial result for abstract LLL was established by Erdős and Lovász, and the first asymmetric version (Lemma 6) was introduced later. In the following, we will introduce a series of LLL variants that offer trade-offs between bounds and complexity.
Lemma 6
(Quantitative Lovász Local Lemma) Given a dependency graph $G=([n],E)$, and a probability vector $\mathbf{p}=(P(A_{1}),\ldots,P(A_{n}))\in(0,1)^{n}$, if there exist real numbers $x_{1},\ldots,x_{n}\in(0,1)$ such that
$P(A_{i})\leq x_{i}\prod_{j\in\mathcal{N}(i)}(1-x_{j})$ (F.1)
for any $i\in[n]$, then
$\Pr(\cap_{A\in\mathcal{A}}\bar{A})>\prod_{j\in[n]}(1-x_{j})$ (F.2)
By utilizing the relationship between the Lovász Local Lemma and the partition function of the hard-core lattice gas on graphs, an improved version of the Lovász Local Lemma was obtained in [2].
Lemma 7
(Quantitative Cluster Expansion Lovász Local Lemma) Given a dependency graph $G=([n],E)$, and a probability vector $\mathbf{p}=(P(A_{1}),\ldots,P(A_{n}))\in(0,1)^{n}$. Suppose that a weighting function $\tilde{\mu}:\mathcal{A}\rightarrow[0,\infty)$ satisfies
$\tilde{\mu}(A_{i})\geq P(A_{i})\times\sum_{I\in Ind(\mathcal{N}^{+}(A_{i}))}\prod_{A_{j}\in I}\mu(A_{j})$ (F.3)
for any $i\in[n]$, where $Ind(G)$ is the collection of independent sets of $G$, then
$\Pr(\cap_{A\in\mathcal{A}}\bar{A})>\prod_{j\in[n]}(1-P(A_{j}))^{\varphi_{\tilde{\mu}}(A_{j})}$ (F.4)
with
$\varphi_{\tilde{\mu}}(A_{j})=\sum_{I\in Ind(\mathcal{N}(A_{j}))}\prod_{A_{k}\in I}\mu(A_{k})$ (F.5)
The original LLL (Lemma 1) only uses minimalistic local structural information. In [15], a hierarchy of LLLs is presented, which are increasingly complex and utilize an increasing amount of local information in a non-trivial way. In the limit, they yield Shearer’s condition, which uses all the global information regarding the structure of the dependency graph.
Lemma 8
(Quantitative Clique Lovász Local Lemma) Let $\mathcal{A}=\{A_{1},A_{2},\ldots,A_{n}\}$ be a set of events with dependency graph $G$ and let $\mathcal{K}=\{K_{1},K_{2},\ldots,K_{m}\}$ be a set of cliques in $G$ covering all the edges (not necessarily disjointly). If there exist a set of vectors $\{\textbf{x}_{1},\textbf{x}_{2},\ldots,\textbf{x}_{m}\}$ from $(0,1)^{n}$ such that following conditions are satisfied:
  1. 1.
    $\forall v\in[m]:\sum_{i\in K_{v}}x_{i,v}<1$
  2. 2.
    $\forall i\in[n],\forall v$ such that $i\in K_{v}$:
    $P(A_{i})\leq x_{i,v}\prod_{u\neq v:i\in K_{u}}(1-\sum_{j\in K_{u}\backslash\{i\}}x_{j,u})$ (F.6)
then:
  1. 1.
    $\Pr(\bigcap_{i\in[n]}\bar{A_{i}})\geq\prod_{v\in[m]}(1-\sum_{i\in K_{v}}x_{i,v})>0$
  2. 2.
    In the variable framework, the running time of the algorithm of Moser and Tardos is at most:
    $\sum_{i\in[n]}min_{v:i\in K_{v}}\frac{x_{i,v}}{1-\sum_{j\in K_{v}}x_{j,v}}$ (F.7)
Algorithm 1 The Moser-Tardos (MT) Algorithm
0: Sample space: $\mathbf{a}=(a_{0},a_{1},\ldots,a_{m_{t}})$.Independent r.v. taking values in $\mathbf{a}$: $\chi=\{X_{1},...,X_{\gamma\kappa}\}$.Set of (ordered) events: $\mathcal{H}=\{H_{1},...,H_{k}\}.$Probability distribution: $\mathbf{p}=(p_{0},p_{1},\ldots,p_{m_{t}})$.
0: Assignment $\alpha=(r_{1},...,r_{\gamma\kappa})\in\mathbf{a}^{\gamma\kappa}$ to r.v. $\chi$ s.t. $\cap_{H_{i}\in\mathcal{H}}\bar{H_{i}}=TRUE$.
1: Sample the variables $X_{i},~i\in[\gamma\kappa]$, and let $\alpha$ be the resulting assignment.
2:while there exists a bad event in $\mathcal{H}$ that occurs under the current assignment, let $H_{j}$ be the least indexed such event do
3:  RESAMPLE($H_{j}$)
4:end while
5: Output current assignment $\alpha$.
RESAMPLE($H_{j}$)
1: Resample the variables in sc($H_{j}$).
2:while there is a least indexed bad event $H_{l}$, such that sc($H_{j}$) $\cap$ sc($H_{l}$) $\neq\emptyset$, occurring under the current assignment, do
3:  RESAMPLE($H_{l}$)
4:end while